OpenStudy (anonymous):

How is the process of adding and subtracting radicals similar to that of simplifying expressions with variables?

OpenStudy (amistre64):

do you know the value of a radical?

OpenStudy (amistre64):

the distributive property works regardless of what the known or unknown values are.

OpenStudy (thejax):

The two main operations you can perform when adding and subtracting radicals are: simplifying radicals for example \[√27=√3×9=√3*√9√= 3√3\] and, once you have done this, gathering similar terms, for example \[√2+3√2=(1+3)√2 =4√2\] Simplification is similar to the process in algebra of removing common factors from an expression, and gathering similar terms is very similar to gathering or combining terms in algebra.

OpenStudy (anonymous):

the radical is treated exactly like a variable in a polynomial expression, following all of the same rules with one exception: if you square a radical it removes the radical over the term inside of it.

OpenStudy (amistre64):

squaring a radical does not remove the radical ...

OpenStudy (amistre64):

\[(\sqrt[3]{5})^2\ne5\]

OpenStudy (anonymous):

ok, ok, technicality. If you square a square root, then it removes the radical. If you cube a cube root, then it removes the radical. Good catch, I wouldn't want someone following that to the letter not understanding what I intended to say. @amistre64